OpenStudy (anonymous):

88 rounded to the nearest tenth

OpenStudy (curry):

0.9

OpenStudy (anonymous):



OpenStudy (anonymous):

thats what the prob looks like

OpenStudy (curry):

16^2-11^2=x^2

OpenStudy (curry):

leace medal if it helped

OpenStudy (anonymous):

a For the following find teh side length x . Round your answer oto the nearest tenth

OpenStudy (curry):

256-121

OpenStudy (curry):

sqrt 135=x

OpenStudy (curry):

11.6

OpenStudy (lalaly):

find x by pythagoras \[16^2=x^2+11^2\] \[x^2=135\]\[x=\sqrt{135}\]\[x=11.618\]\[x=11.6\]

OpenStudy (anonymous):

thanu ummmm soo 11.6 is the final answer what would it be rounded to the nearest tenth??

OpenStudy (curry):

yes

OpenStudy (lalaly):

11.6 is the final answer

OpenStudy (anonymous):

thnks